# Stability-Certified On-Policy Data-Driven LQR via Recursive Learning and Policy Gradient

### v1

@@ Appendix A Preliminaries on averaging theory for two-time-scale systems
$\leq\gamma\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|\!$
$\hskip 28.45274pt\cdot\left\|G(\tilde{K}^{\textsc{av}}_{t}\!+\!K^{\star},\tilde{\theta}^{\textsc{av}}_{t}\!+\!\theta^{\star})\!-\!G(\tilde{K}^{\textsc{av}}_{t}\!+\!K^{\star},\theta^{\star})\right\|$
@@
$\hskip 11.38092pt+\tfrac{\gamma^{2}\beta_{1}}{2}\left\|G(\tilde{K}^{\textsc{av}}_{t}\!+\!K^{\star},\tilde{\theta}^{\textsc{av}}_{t}\!+\!\theta^{\star})\right\|^{2}$
$\hskip 11.38092pt+\tfrac{\gamma^{2}\beta_{1}}{2}\left\|G(\tilde{K}^{\textsc{av}}_{t}\!+\!K^{\star},\theta^{\star})\right\|^{2}$
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\gamma\beta_{2}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|$
@@
$\hskip 11.38092pt+\tfrac{\gamma^{2}\beta_{1}}{2}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\tilde{\theta}^{\textsc{av}}_{t}+\theta^{\star})\pm G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|^{2}$
@@
$\hskip 11.38092pt+\tfrac{\gamma^{2}\beta_{1}}{2}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|^{2}$
@@
$\stackrel{{\scriptstyle(b)}}{{\leq}}\gamma\beta_{2}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|$
$\hskip 11.38092pt+\gamma^{2}\beta_{1}\beta_{2}\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2}+\gamma^{2}\beta_{1}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|^{2}$
@@
$\hskip 11.38092pt+\tfrac{\gamma^{2}\beta_{1}}{2}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|^{2},$ (D.5)
@@
where in $(a)$ we exploited the Lipschitz continuity expressed on (D.3b) and added ±$\pm G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})$ inside the norm of the second term, while in $(b)$ we exploited again the Lipschitz continuity and a standard property of the square norm [^7] Given any $v_{1},v_{2}\!\in\!\mathbb{R}^{n}$, it holds $\left\|v_{1}\!-\!v_{2}\right\|^{2}\!\leq\!2\left\|v_{1}\right\|^{2}\!+\!2\left\|v_{2}\right\|^{2}$.. Plugging the bound in (D.5) into (D.4) and restricting $\kappa\in(0,1)$, we get
@@
$\Delta V(\tilde{K}^{\textsc{av}}_{t},\tilde{\theta}^{\textsc{av}}_{t})$ $\leq-\gamma\kappa\left(1-\gamma\tfrac{3\beta_{1}}{2}\right)\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|^{2}$
@@
$\hskip 11.38092pt+\gamma\kappa\beta_{2}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|$
@@
$\hskip 11.38092pt-\gamma\left(1-\gamma\tfrac{1+\beta_{1}\beta_{2}^{2}}{2}\right)\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2},$ (D.6)
@@
where we used $\kappa\beta_{1}\beta_{2}^{2}\leq\beta_{1}\beta_{2}^{2}$. Let us arbitrarily choose $\nu_{1},\nu_{2}\in(0,1)$. Then, for all $\gamma\in(0,\bar{\gamma}^{\textsc{av}})$ with $\bar{\gamma}^{\textsc{av}}:=\min\left\{1,\frac{2\nu_{1}}{3\beta_{1}},\frac{2\nu_{2}}{1+\beta_{1}\beta_{2}^{2}}\right\}$, we further bound (D.6) as
$\Delta V(\tilde{K}^{\textsc{av}}_{t},\tilde{\theta}^{\textsc{av}}_{t})$
$\leq-\gamma\kappa\nu_{1}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|^{2}$
@@
$\hskip 11.38092pt+\gamma\kappa\beta_{2}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|-\gamma\nu_{2}\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2}$
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}-\gamma\begin{bmatrix}\left\|G(\tilde{K}^{\textsc{av}}_{t}\!+\!K^{\star},\theta^{\star})\right\|\\ \left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|\end{bmatrix}^{\top}\!\!\!U(\kappa)\!\begin{bmatrix}\left\|G(\tilde{K}^{\textsc{av}}_{t}\!+\!K^{\star},\theta^{\star})\right\|\\ \left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|\end{bmatrix}\!,$ (D.11)
@@
$U(\kappa):=\begin{bmatrix}\kappa\nu_{1}&-\tfrac{\kappa\beta_{2}}{2}\\ -\tfrac{\kappa\beta_{2}}{2}&\nu_{2}\end{bmatrix}.$
In light of the Sylvester criterion, the matrix $U(\kappa)$ is positive definite if and only if its determinant is positive. Hence, we further restrict $\kappa\in(0,\bar{\kappa})$, with $\bar{\kappa}:=\min\{\beta_{2}/(4\nu_{1}\nu_{2}),1\}$. Let $\eta>0$ be the smallest eigenvalue of $U(\kappa)$, then (D.11) can be bounded as
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}-\gamma\tfrac{\eta}{\mu}\left(J(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})-J(K^{\star},\theta^{\star})\right)$
$\hskip 11.38092pt-\gamma\eta\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}-\gamma\eta\min\left\{\tfrac{1}{\mu\kappa},1\right\}V(\tilde{\theta}^{\textsc{av}}_{t},\tilde{K}^{\textsc{av}}_{t}),\!$ (D.12)
where $(a)$ follows from the gradient dominance of $J$ (cf. (D.1)), while $(b)$ recovers the formulation of $V$ (cf. (D.2)) by negleting a negative term. Being the right-hand side of (D.12) always non-positive, it holds
$V(\tilde{K}^{\textsc{av}}_{{t+1}},\tilde{\theta}^{\textsc{av}}_{{t+1}})$ $\leq V(\tilde{K}^{\textsc{av}}_{t},\tilde{\theta}^{\textsc{av}}_{t})$
$\stackrel{{\scriptstyle(a)}}{{\leq}}J(\tilde{K}^{\textsc{av}}_{t}\!+\!K^{\star},\theta^{\star})\!-\!J(K^{\star},\theta^{\star})\!+\!\tfrac{1}{2}\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2}\!\!,$
@@
Since system (40) is initialized into $\Omega_{c_{0}}$, then its trajectories satisfy (D.12). Hence, the exponential stability of the origin for system (40) is implied (cf. [56, Theorem 13.2]) and the proof follows.

### v2

@@ 6 Conclusions
$\leq\gamma\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|\!\left\|G(\tilde{K}^{\textsc{av}}_{t}\!+\!K^{\star},\tilde{\theta}^{\textsc{av}}_{t}\!+\!\theta^{\star})\!-\!G(\tilde{K}^{\textsc{av}}_{t}\!+\!K^{\star},\theta^{\star})\right\|$
@@
$\hskip 11.38092pt+\tfrac{\gamma^{2}\beta_{3}}{2}\left\|G(\tilde{K}^{\textsc{av}}_{t}\!+\!K^{\star},\tilde{\theta}^{\textsc{av}}_{t}\!+\!\theta^{\star})\right\|^{2}+\tfrac{\gamma^{2}\beta_{3}}{2}\left\|G(\tilde{K}^{\textsc{av}}_{t}\!+\!K^{\star},\theta^{\star})\right\|^{2}$
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\gamma\beta_{4}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|$
@@
$\hskip 11.38092pt+\tfrac{\gamma^{2}\beta_{3}}{2}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\tilde{\theta}^{\textsc{av}}_{t}+\theta^{\star})\pm G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|^{2}$
@@
$\hskip 11.38092pt+\tfrac{\gamma^{2}\beta_{3}}{2}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|^{2}$
@@
$\stackrel{{\scriptstyle(b)}}{{\leq}}\gamma\beta_{4}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|+\gamma^{2}\beta_{3}\beta_{4}\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2}+\gamma^{2}\beta_{3}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|^{2}$
@@
$\hskip 11.38092pt+\tfrac{\gamma^{2}\beta_{3}}{2}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|^{2},$ (C.7)
@@
where in $(a)$ we exploited the Lipschitz continuity expressed on (C.5b) and added ±$\pm G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})$ inside the norm of the second term, while in $(b)$ we exploited again the Lipschitz continuity and a standard property of the square norm [^7] Given any $v_{1},v_{2}\!\in\!\mathbb{R}^{n}$, it holds $\left\|v_{1}\!-\!v_{2}\right\|^{2}\!\leq\!2\left\|v_{1}\right\|^{2}\!+\!2\left\|v_{2}\right\|^{2}$.. Plugging the bound in (C.7) into (C.6) and restricting $\kappa\in(0,1)$, we get
@@
$\Delta V(\tilde{K}^{\textsc{av}}_{t},\tilde{\theta}^{\textsc{av}}_{t})$ $\leq-\gamma\kappa\left(1-\gamma\tfrac{3\beta_{3}}{2}\right)\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|^{2}$
@@
$\hskip 11.38092pt+\gamma\kappa\beta_{4}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|$
@@
$\hskip 11.38092pt-\gamma\left(1-\gamma\tfrac{1+\beta_{3}\beta_{4}^{2}}{2}\right)\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2},$ (C.8)
@@
where we used $\kappa\beta_{3}\beta_{4}^{2}\leq\beta_{3}\beta_{4}^{2}$. Let us arbitrarily choose $\nu_{1},\nu_{2}\in(0,1)$. Then, for all $\gamma\in(0,\bar{\gamma}^{\textsc{av}})$ with $\bar{\gamma}^{\textsc{av}}:=\min\left\{1,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\gamma}_{0}},\frac{2\nu_{1}}{3\beta_{3}},\frac{2\nu_{2}}{1+\beta_{3}\beta_{4}^{2}}\right\}$, we further bound (C.8) as
@@
$\Delta V(\tilde{K}^{\textsc{av}}_{t},\tilde{\theta}^{\textsc{av}}_{t})$ $\leq-\gamma\kappa\nu_{1}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|^{2}+\gamma\kappa\beta_{4}\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|$
$\hskip 11.38092pt-\gamma\nu_{2}\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2}$
@@
=$\!\stackrel{{\scriptstyle(a)}}{{=}}\!-\gamma\begin{bmatrix}\left\|G(\tilde{K}^{\textsc{av}}_{t}\!+\!K^{\star},\theta^{\star})\right\|\\ \left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|\end{bmatrix}^{\top}\!\!\!U(\kappa)\!\!\!\begin{bmatrix}\left\|G(\tilde{K}^{\textsc{av}}_{t}\!+\!K^{\star},\theta^{\star})\right\|\\ \left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|\end{bmatrix}\!\!,\!\!$ (C.9)
@@
$U(\kappa):=\begin{bmatrix}\kappa\nu_{1}&-\tfrac{\kappa\beta_{4}}{2}\\ -\tfrac{\kappa\beta_{4}}{2}&\nu_{2}\end{bmatrix}.$
In light of the Sylvester criterion, the matrix $U(\kappa)$ is positive definite if and only if its determinant is positive. Hence, we further restrict $\kappa\in(0,\bar{\kappa})$, with $\bar{\kappa}:=\min\{\beta_{4}/(4\nu_{1}\nu_{2}),1\}$. Let $\eta>0$ be the smallest eigenvalue of $U(\kappa)$, then (C.9) can be bounded as
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}-\gamma\tfrac{\eta}{\mu}\left(J(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})-J(K^{\star},\theta^{\star})\right)-\gamma\eta\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}-\gamma\eta\min\left\{\tfrac{1}{\mu\kappa},1\right\}V(\tilde{\theta}^{\textsc{av}}_{t},\tilde{K}^{\textsc{av}}_{t}),\!$ (C.10)
where $(a)$ follows from the gradient dominance of $J$ (cf. (C.1)), while $(b)$ recovers the formulation of $V$ (cf. (C.2)) by negleting a negative term. Being the right-hand side of (C.10) always non-positive, it holds
$V(\tilde{K}^{\textsc{av}}_{{t+1}},\tilde{\theta}^{\textsc{av}}_{{t+1}})$ $\leq V(\tilde{K}^{\textsc{av}}_{t},\tilde{\theta}^{\textsc{av}}_{t})\stackrel{{\scriptstyle(a)}}{{\leq}}J(\tilde{K}^{\textsc{av}}_{t}\!+\!K^{\star},\theta^{\star})\!-\!J(K^{\star},\theta^{\star})\!+\!\tfrac{1}{2}\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2}\!\!,$
@@
Since system (36) is initialized into $\Omega_{c_{0}}$, then its trajectories satisfy (C.10). Hence, the exponential stability of the origin for system (36) is implied (cf. [57, Theorem 13.2]) and the proof follows.